OpenStudy (anonymous):

would 5sqrtx^25 be x^5?

jimthompson5910 (jim_thompson5910):

Yes, \[\large \sqrt[5]{x^{25}}=x^5\]

jimthompson5910 (jim_thompson5910):

This is because \[\large \sqrt[5]{x^{25}}=(x^{25})^{\frac{1}{5}}=x^{\frac{25}{5}}=x^5\]

OpenStudy (anonymous):

thank you! I just may be finally getting it! :)

jimthompson5910 (jim_thompson5910):

that's great

OpenStudy (anonymous):

even without using rational exponents you can say "5 goes in to 25 5 times"

OpenStudy (anonymous):

yes that's how I came up with it

OpenStudy (anonymous):

I dont think so. coz 5sqrt(x^25) = 5(x^(25/2)) and there is not a clear way (for me) to simplify it